### v1

@@ 3 Fairness Concerns in Healthcare
This section delves into various fairness concerns that arise during the application of AI in healthcare. Initially, it examines the sources of bias in AI healthcare systems, including data bias, algorithm bias, explicit bias, implicit bias, and selection bias. Furthermore, it will discuss the potential consequences of these biases, such as misdiagnosis, inequitable healthcare outcomes, loss of trust in healthcare systems, legal and ethical implications, resource misallocation, and stifling innovation. By delving into these topics, this section aims to underscore the critical need to address biases to ensure AI technologies contribute to fair and equitable healthcare for all.
@@ 3.1 Sources of bias
Data Bias: Data bias in AI healthcare systems can severely compromise their effectiveness and fairness. For instance, a study by Obermeyer et al. 12 found that a healthcare management algorithm disproportionately favored white patients over black patients with similar health conditions due to reliance on historical healthcare cost data, which reflected existing socioeconomic disparities. Further studies reinforce the pervasive nature of data bias in healthcare AI. For example, Rajkomar and his team found that the models trained on patient datasets from one geographical region often failed to generalize to populations outside that region, indicating a geographic data bias that limits the practical utility of AI systems in diverse settings 4. Additionally, the AI system used for diagnosing diabetic retinopathy performed well in tests but showed decreased accuracy in clinical settings, especially among minority populations not well represented in the training data, underscoring the impact of demographic data biases 116.
Algorithm Bias: Algorithmic bias in AI healthcare occurs when AI systems develop systematic errors that disadvantage certain patient groups. For example, AI models used to diagnose pneumonia displayed biases against specific racial groups, failing to capture important socio-economic and demographic factors despite extensive training on hospital data. This led to inappropriate treatment recommendations 4, 13, 117. Another study in dermatology faced the same issue. The CNNs used for detecting skin cancer were less effective for individuals with darker skin. This discrepancy arose because the training data predominantly consisted of images featuring lighter skin tones, which does not reflect the diversity of global skin types. Consequently, this training limitation can lead to poorer health outcomes for underrepresented groups 14.
Explicit Bias: Explicit bias in AI systems within healthcare refers to observable prejudices that directly influence these systems’ decision-making processes. These biases are often a result of human prejudices being encoded in to AI through biased data or algorithms, leading to discriminatory practices and unequal treatment of patients 118. One example of explicit bias in healthcare AI is the underdiagnosis of certain demographic groups by AI models. A study investigating the fairness of vision language foundation models in medical imaging found that these AI systems consistently underdiagnosed marginalized groups, such as black female patients, across a range of pathologies 119. This explicit bias in diagnostic AI tools can lead to delayed or inadequate medical care for specific patient subgroups, exacerbating healthcare disparities. Underlying Clinical Decision Support systems, explicit bias may be due to the selection of data and algorithm design, leading to biased outcomes. For example, the decreased likelihood of receiving diagnostic imaging for nonwhite groups compared to their white counterparts when controlling for patient and facility level factors is embedded in clinical data 120. Accordingly, models predicting imaging use may reflect the human prejudices around who deserves and receives diagnostics 121. Moreover, if an AI system is trained predominantly on data from one ethnic group or gender, it may not perform accurately for other ethnicities or gender. The use of biased AI in healthcare can result in systematic discrimination, where certain groups received suboptimal care due to the AI’s prejudiced programming. This can manifest in various ways, such as through biased diagnostic tools, treatment recommendations, or predictive analytics that unfairly disadvantage certain groups.
Implicit Bias: Implicit bias in AI systems in healthcare refers to the unconscious prejudices that can influence decision-making processes without explicit acknowledgment. These biases often reflect societal stereotypes and may not align with the conscious values or beliefs of the developers or users of the AI system. Implicit biases can alter perceptions and judgments, leading to disparities in healthcare delivery and outcomes 119. In healthcare settings, implicit biases can manifest in various ways. For example, a study on obstetric care providers found that implicit biases favored French women over African migrant women, potentially influencing medical decisions and leading to healthcare disparities 122. Similarly, healthcare professionals have been shown to exhibit implicit biases based on ethnicity, which could affect the likelihood of patients being given another opportunity to attend an outpatient clinic after missing an appointment 123. Another example is the use of AI in diagnostic tools, where implicit biases can result in differential accuracy rates across racial or ethnic groups. This can lead to underdiagnosis or misdiagnosis for certain populations, contributing to unequal treatment and outcomes 124. Implicit biases can also affect the development of clinical decision support systems. This is because the data used to train AI might not fully reflect the diversity of the patient population, which could lead to recommendations and treatment plans that aren’t based on facts 125, 126. Furthermore, implicit biases can impact the patient-provider relationship, where unconscious prejudices may influence the level of care and communication provided to patients of different racial or ethnic backgrounds, potentially affecting their health outcomes 127.
Selection Bias: Selection bias occurs when the training data for an AI model is not representative of the broader population or the context in which the model will be applied. This leads to AI systems performing well on training data but poorly in real-world settings, particularly for underrepresented groups 128. The development of AI models for disease prediction or diagnosis, primarily trained on data from specific demographic groups like a particular race, gender, or age group, is one example of selection bias in healthcare settings. If an AI model for diagnosing skin cancer is trained primarily on images of lighter-skinned individuals, it may perform poorly when diagnosing conditions on darker skin, leading to higher misdiagnosis rates for individuals with darker skin tones 125. AI models that predict patient outcomes or treatment efficacy also exhibit selection bias. If the historical data used to train these models predominantly come from patients who received a certain level of care or who have specific socioeconomic backgrounds, the AI’s predictions may not accurately reflect outcomes for patients outside of these groups 128. This can lead to disparities in healthcare recommendations and interventions offered to patients from different socioeconomic or racial backgrounds. Furthermore, selection bias can also manifest in AI-driven tools for patient triage and resource allocation. If the data used to train these models does not adequately account for the diversity of patient presentations and conditions, the AI system may prioritize certain patient groups over others, inadvertently exacerbating existing healthcare disparities.
@@ 3.2 Potential consequences of biases
Misdiagnosis and Inequitable Healthcare Outcomes: AI integration in healthcare brings significant advancements but also risks of misdiagnosis and inequitable outcomes due to biased algorithms. AI systems often rely on non-representative data, leading to biased decision-making. For instance, Obermeyer highlighted that an algorithm used in healthcare disproportionately favored white patients over black patients because it used healthcare costs as a proxy for healthcare needs, indirectly embedding racial biases in its predictions 12. Moreover, these biases in AI can exacerbate existing healthcare disparities. Rajkomar and others discussed how AI applications, if not carefully designed and monitored, could inherit and amplify socioeconomic and racial disparities. This is particularly problematic in diagnostics, where AI systems are trained predominantly on data from specific demographic groups, potentially leading to poorer diagnostic accuracy for underrepresented groups 4. This was evident in a study by Adamson and Smith, which found that dermatology AI systems demonstrated lower accuracy rates in skin lesion diagnosis for dark-skinned individuals compared to those with lighter skin 14. Such biased AI not only risks misdiagnosis but also contributes to inequitable healthcare outcomes by potentially steering healthcare resources away from those who may need them most. Vayena and his team emphasized the ethical imperative to ensure that AI tools in healthcare are developed with consideration for fairness and equity, advocating for diverse and inclusive data sets and algorithmic transparency to mitigate these biases 129.
@@
Loss of trust in Healthcare Systems: Recent studies have echoed the concern that the deployment of biased AI in healthcare could significantly undermine public trust in medical systems. Trust is foundational to the patient-provider relationship and is crucial for the effective delivery of healthcare services 130. When AI tools exhibit bias, whether in diagnosis, treatment recommendations, or patient management, they can lead to misdirected care, fostering distrust among patients, particularly in marginalized communities 4. Kherbache and his team pointed out that when patients perceive AI-driven processes as opaque or unfair, their trust in the overall healthcare system may decline 131. Incidents of AI failures that receive public attention can exacerbate this erosion of trust, leading patients to question the reliability and ethics of using AI in medical decision-making. Researchers like Kerasidou argue that trust is not only about the accuracy of AI but also its alignment with ethical principles that govern healthcare, such as beneficence and non-maleficence 132. Furthermore, the lack of transparency in how AI models make decisions can be a major barrier to trust. Blease’s study 133 suggests that without clear communication about how AI tools contribute to healthcare decisions, patients may become skeptical of diagnoses and treatments, fearing that their personal healthcare data could be misused or misunderstood. This potential trust deficit could have severe implications, not just for individual health outcomes but also for public health at large, as mistrust in healthcare systems can lead to lower rates of healthcare utilization, vaccine hesitancy, and poor adherence to medical advice 134. Veinot highlights the need for healthcare systems to maintain high standards of accountability and transparency as they integrate AI technologies to mitigate such risks 135.
@@
Legal and Ethical Implications: The use of biased AI in healthcare not only poses clinical risks but also entails significant legal and ethical implications. Deploying biased AI systems could lead to legal breaches of anti-discrimination laws. In the United States, for example, the Civil Rights Act and the Americans with Disabilities Act set legal standards that could be violated by biased AI algorithms, which fail to provide equitable care across different patient demographics 136. Ethically, biased AI conflicts with the fundamental medical ethics principles of justice and non-maleficence, demanding fairness and avoidance of harm, respectively 137. Moreover, biased algorithms in healthcare could potentially expose medical practices and institutions to litigation related to malpractice or negligence, especially if these algorithms contribute to substandard care outcomes 138. For instance, if an AI system were to consistently provide inferior diagnostic support for certain racial groups, this could be viewed as a form of systemic negligence or malpractice 129. Additionally, the ethical implications extend to the breach of patient trust and the compromise of patient autonomy. Ethical medical practice relies heavily on the principles of informed consent and respect for patients’ autonomy—principles that are challenged by opaque AI systems that do not make their decision-making processes or inherent biases clear to patients or practitioners 133. These potential legal and ethical failures underscore the necessity for rigorous oversight and transparent development processes for AI in healthcare, aiming to ensure that these technologies adhere to both existing legal frameworks and ethical standards of practice.
@@
Resource Misallocation: The deployment of biased AI in healthcare settings can lead to resource misallocation, a critical issue that impacts both the efficiency and fairness of medical services. Biased algorithms may misdirect resources by prioritizing certain groups over others based on flawed data inputs or biased training procedures. For example, a study by Obermeyer demonstrated how an algorithm used for managing healthcare resources inadvertently favored healthier white patients over sicker black patients due to biased data inputs that did not accurately reflect patient needs 12. This misallocation can exacerbate existing healthcare disparities by diverting necessary medical attention and resources away from those who are most in need. As Chen pointed out, such disparities are not just a matter of clinical outcomes but are deeply tied to social and economic inequalities that AI tools can inadvertently perpetuate 13. Furthermore, biased AI can influence the allocation of resources within healthcare facilities, potentially resulting in inefficiencies that strain healthcare systems. This includes misallocating medical staff, diagnostic tools, and hospital beds, which can degrade the quality of care delivered while increasing wait times and healthcare costs 4. Resource misallocation also raises ethical questions about fairness and equity in healthcare provisioning. It challenges the ethical principle of justice, which demands that healthcare resources be distributed based on need rather than biased algorithms 137. This ethical breach can lead to further mistrust and reluctance among underserved populations to engage with healthcare systems, perpetuating a cycle of disadvantage.
@@
Stifling Innovation: The presence of bias in AI systems used in healthcare not only affects the accuracy and fairness of medical services but can also stifle innovation. When AI models are made using biased data, they might not accurately reflect the different needs of the population. This could make these innovations less useful and applicable to a wider range of demographic groups. Liang talks about how biased training datasets in AI development make it harder for these systems to work with different types of data. This could stymie innovation by preventing it from solving larger or more complex health problems that affect many people. Moreover, reliance on biased AI can deter investment in developing new technologies that are inclusive and equitable. Potential investors may be cautious about funding projects that might not meet regulatory standards for fairness or could lead to public backlash — a concern highlighted by Corbett-Davies and Goel, who note the legal and social implications of deploying biased AI 139. In addition, the perpetuation of bias in AI could solidify existing disparities in healthcare innovation. As Benjamin notes, if innovation is directed predominantly at solving problems specific to well-represented groups, less common but equally pressing issues in underrepresented groups are neglected, thereby limiting the scope and impact of technological advancements in healthcare 140. The studies of Rajkomar 4, Hardt 141, and Howell 142 made things even more complicated by pointing out that biased algorithms can make it harder to find new treatments that work for everyone. This feeds a cycle of innovation that helps groups that are already overrepresented in data sets more than others. This restricted focus on innovation not only affects the equity of healthcare delivery but also limits the development of truly innovative, comprehensive healthcare solutions. These factors combined suggest that biased AI not only hinders the progression of medical technologies but also potentially locks the healthcare sector into a cycle of uneven innovation where only the needs of the majority are systematically addressed.

### v2

@@ 2 Fairness Concerns in Healthcare
This section explores the fairness concerns associated with the use of AI in healthcare through three key areas. It begins by examining the sources of bias that can emerge at various stages of the machine learning pipeline. Next, it discusses the potential consequences of these biases, including inequitable healthcare outcomes, loss of trust, and ethical and legal challenges. Finally, it introduces fairness metrics and trade-offs, which provide a framework for addressing bias while balancing the inherent tensions between fairness and performance in healthcare applications. By delving into these topics, this section aims to underscore the critical need to address biases to ensure AI technologies contribute to fair and equitable healthcare for all.
@@ 2.1 Sources of Bias
Bias in AI healthcare systems can arise at multiple stages of the machine learning (ML) pipeline, necessitating phase-specific mitigation strategies. To address this comprehensively, researchers have developed three primary approaches, each targeting a distinct phase of the pipeline: (i) pre-processing, which focuses on reducing bias in the input data before model training [59]; (ii) in-processing, which applies fairness constraints during the training phase [60]; and (iii) post-processing, which modifies model outputs to enhance fairness [61]. We adopt this three-stage framework to capture the end-to-end process of data collection, model development, and decision-making in AI healthcare systems. This structure, widely used in the fairness literature, supports systematic bias analysis and enables targeted detection and mitigation strategies at each stage [62, 59, 60, 61, 63, 64]. The division also facilitates clearer reasoning about when and how fairness interventions can be most effectively applied throughout the ML pipeline.
A critical first step in designing effective interventions is to identify the underlying sources of bias at each stage of the ML pipeline. This section outlines those sources, with a summary provided in Table 1 and detailed discussions presented in the following sections.
Table 1: Summary of biases across each stage of the machine learning (ML) pipeline
Stage Bias Type References Description
Pre-processing Selection Bias [65, 66] Arises from how the data is chosen or sampled for training, leading to unrepresentative datasets.
Measurement Bias [62, 65, 67] Results from the ways features are chosen, utilized, or measured, leading to the systematic over- or under-representation of certain groups or variables in the data.
Representation Bias [65, 14] Occurs when datasets fail to adequately capture the diversity of the populations they aim to serve.
Explicit Bias [62] Involves intentional or overt decisions that result in discriminatory patterns in data preparation or selection.
Implicit Bias [68] Stems from unintentional or unconscious decisions that lead to skewed datasets, frequently influenced by systemic inequities.
In-processing Algorithmic Bias [69, 70] Arises when algorithms magnify existing biases in the training data or when they are inherently biased due to their design.
Explicit Bias [62] Involves deliberate manipulations within the model’s algorithms or training process that introduce or amplify disparities.
Implicit Bias [71] Arises from model design decisions that unintentionally embed existing societal inequities, often due to insufficient awareness or oversight.
Post-processing Evaluation Bias [72] Arises when performance assessments fail to account for disparities across demographic groups, leading to inequities in clinical outcomes.
Explicit Bias [62] Involves intentional manipulation of outputs or decision thresholds that disproportionately affect specific groups.
Implicit Bias [68] Occurs when decision-making criteria unintentionally create disparities, often because important fairness factors are overlooked.
2.1.1 Pre-Processing Stage
In the pre-processing stage, bias can originate from data collection, preparation, or representation before modeling begins. Because both the performance and fairness of an AI system depend on the quality of its data, bias at this stage can have widespread consequences. One common form of bias is selection bias, which arises from how the data is chosen or sampled for training, leading to unrepresentative datasets [65]. For instance, selection bias occurs when an AI model is trained predominantly on data from one demographic (e.g., white patients), while underrepresented groups, such as Black patients, have limited representation or inaccessible health records. This can lead to false negatives, fewer follow-up scans, and undiagnosed conditions, ultimately worsening health inequities for disadvantaged populations [66].
Another form of bias is measurement bias, which results from the ways features are chosen, utilized, or measured, leading to the systematic over- or under-representation of certain groups or variables in the data [62, 65]. An example is the use of pulse oximeters, devices that estimate blood oxygen levels by emitting light through the skin. Research indicates that pulse oximeters often provide less accurate readings for individuals with darker skin tones, frequently overestimating their oxygen saturation levels [67]. This discrepancy can result in the underdiagnosis or delayed treatment of hypoxemia in these populations, rendering the models unreliable or biased in their predictions. Similarly, representation bias occurs when datasets fail to adequately capture the diversity of the populations they aim to serve [65]. For example, dermatological AI systems trained primarily on images of lighter skin tones often underperform when diagnosing conditions on darker skin [14]. This lack of representation can lead to inequitable healthcare outcomes and limit the generalizability of the AI system.
Within pre-processing, biases can also appear explicitly or implicitly. Explicit bias involves intentional or overt decisions that result in discriminatory patterns in data preparation or selection [62]. For instance, consider a dataset for training an AI model to predict heart disease. If the dataset is curated to include only data from male patients and excludes female patients entirely, this introduces explicit bias. This exclusion is deliberate and creates a model that cannot generalize to female patients, potentially leading to disparities in healthcare outcomes. Implicit bias, on the other hand, often stems from unintentional or unconscious decisions that lead to skewed datasets, frequently influenced by systemic inequities [68]. For example, using historical healthcare data that disproportionately underrepresented women in clinical trials introduces an implicit bias into the AI model. The unconscious reliance on historically biased datasets perpetuates existing inequalities, such as the underdiagnosis of heart attacks in women.
2.1.2 In-Processing Stage
In the in-processing stage, biases can emerge during model training or can be introduced through algorithmic choices, significantly impacting an AI system’s outputs. Although the data itself can carry biases, the algorithms that learn from this data can also amplify or introduce a form of bias often referred to as algorithmic bias. It arises when algorithms magnify existing biases in the training data or when they are inherently biased due to their design [69]. For example, the Vaginal Birth After Cesarean (VBAC) calculator included race-based correction factors that systematically assigned lower success probabilities to African American and Hispanic women, discouraging VBAC attempts for these groups without robust scientific justification [70]. Such a design flaw not only reflected existing disparities in maternal healthcare but also exacerbated them, highlighting how biased algorithms can influence decision-making in critical areas.
Explicit and implicit biases can also arise during in-processing. Explicit bias involves deliberate manipulations within the model’s algorithms or training process that introduce or amplify disparities [62]. For instance, intentionally adjusting the algorithm to prioritize predictions for one demographic group over another (e.g., optimizing cancer detection accuracy only for men) constitutes explicit bias. Such direct alterations can skew the model’s performance metrics, favoring specific groups while disadvantaging others, resulting in ethical and clinical concerns. Implicit bias, however, arises from model design decisions that unintentionally embed existing societal inequities, often due to insufficient awareness or oversight [71]. For example, training a diabetes prediction model using an optimization metric that inadvertently prioritizes accuracy over fairness for underserved populations can perpetuate health disparities. These unconscious design choices may cause the model to fail to serve vulnerable groups adequately.
2.1.3 Post-Processing Stage
Finally, the post-processing stage involves how model outputs are evaluated, interpreted, and deployed in real clinical settings. Evaluation bias arises when performance assessments fail to account for disparities across demographic groups, leading to inequities in clinical outcomes [72]. For example, if an AI model for detecting diabetic retinopathy is evaluated primarily on data from urban hospitals, its performance metrics may overstate its accuracy and fail to reflect reduced performance in rural or underserved populations. This oversight in the evaluation process can lead to biased deployment decisions, disproportionately affecting access to reliable diagnostics in marginalized communities.
As in earlier stages, biases at this point can be explicit or implicit. Explicit bias involves intentional manipulation of outputs or decision thresholds that disproportionately affect specific groups [62]. For instance, adjusting the risk score thresholds for approving diagnostic tests based on race or socioeconomic status. Such actions, when deliberate, can reinforce discriminatory practices in clinical decision-making and resource allocation. Implicit bias, on the other hand, occurs when decision-making criteria unintentionally create disparities, often because important fairness factors are overlooked [68]. Consider an AI system that predicts which patients are at risk of readmission after hospital discharge. During the post-processing stage, the hospital decides to prioritize follow-up care for patients with private insurance, based on an assumption that they are more likely to comply with medical recommendations. However, this approach unintentionally introduces implicit bias, as patients with public insurance or no insurance who may face greater barriers to accessing care are deprioritized. The bias does not stem from a deliberate decision to disadvantage these groups, but rather from an unconscious oversight of fairness considerations in the post-processing criteria.
By examining biases according to their emergence in the ML pipeline, this structured framework facilitates clearer understanding of their origins and supports targeted mitigation strategies. Such an approach underscores the importance of addressing biases early and continuously from data collection and algorithm design to final model deployment and evaluation to ensure equitable and responsible AI-driven healthcare solutions.
@@ 2.2 Potential consequences of biases
Misdiagnosis and Inequitable Healthcare Outcomes: AI integration in healthcare brings significant advancements but also risks of misdiagnosis and inequitable outcomes due to biased algorithms. AI systems often rely on non-representative data, leading to biased decision-making. For instance, Obermeyer highlighted that an algorithm used in healthcare disproportionately favored white patients over black patients because it used healthcare costs as a proxy for healthcare needs, indirectly embedding racial biases in its predictions [12]. Moreover, these biases in AI can exacerbate existing healthcare disparities. Rajkomar and others discussed how AI applications, if not carefully designed and monitored, could inherit and amplify socioeconomic and racial disparities. This is particularly problematic in diagnostics, where AI systems are trained predominantly on data from specific demographic groups, potentially leading to poorer diagnostic accuracy for underrepresented groups [4]. This was evident in a study by Adamson and Smith, which found that dermatology AI systems demonstrated lower accuracy rates in skin lesion diagnosis for dark-skinned individuals compared to those with lighter skin [14]. Such biased AI not only risks misdiagnosis but also contributes to inequitable healthcare outcomes by potentially steering healthcare resources away from those who may need them most. Vayena and his team emphasized the ethical imperative to ensure that AI tools in healthcare are developed with consideration for fairness and equity, advocating for diverse and inclusive data sets and algorithmic transparency to mitigate these biases [73].
@@
Loss of trust in Healthcare Systems: Recent studies have echoed the concern that the deployment of biased AI in healthcare could significantly undermine public trust in medical systems. Trust is foundational to the patient-provider relationship and is crucial for the effective delivery of healthcare services [74]. When AI tools exhibit bias, whether in diagnosis, treatment recommendations, or patient management, they can lead to misdirected care, fostering distrust among patients, particularly in marginalized communities [4]. Kherbache and his team pointed out that when patients perceive AI-driven processes as opaque or unfair, their trust in the overall healthcare system may decline [75]. Incidents of AI failures that receive public attention can exacerbate this erosion of trust, leading patients to question the reliability and ethics of using AI in medical decision-making. Researchers like Kerasidou argue that trust is not only about the accuracy of AI but also its alignment with ethical principles that govern healthcare, such as beneficence and non-maleficence [76]. Furthermore, the lack of transparency in how AI models make decisions can be a major barrier to trust. Blease’s study [77] suggests that without clear communication about how AI tools contribute to healthcare decisions, patients may become skeptical of diagnoses and treatments, fearing that their personal healthcare data could be misused or misunderstood. This potential trust deficit could have severe implications, not just for individual health outcomes but also for public health at large, as mistrust in healthcare systems can lead to lower rates of healthcare utilization, vaccine hesitancy, and poor adherence to medical advice [78]. Veinot highlights the need for healthcare systems to maintain high standards of accountability and transparency as they integrate AI technologies to mitigate such risks [79].
@@
Legal and Ethical Implications: The use of biased AI in healthcare not only poses clinical risks but also entails significant legal and ethical implications. Deploying biased AI systems could lead to legal breaches of anti-discrimination laws. In the United States, for example, the Civil Rights Act and the Americans with Disabilities Act set legal standards that could be violated by biased AI algorithms, which fail to provide equitable care across different patient demographics [80]. Ethically, biased AI conflicts with the fundamental medical ethics principles of justice and non-maleficence, demanding fairness and avoidance of harm, respectively [81]. Moreover, biased algorithms in healthcare could potentially expose medical practices and institutions to litigation related to malpractice or negligence, especially if these algorithms contribute to substandard care outcomes [82]. For instance, if an AI system were to consistently provide inferior diagnostic support for certain racial groups, this could be viewed as a form of systemic negligence or malpractice [73]. Additionally, the ethical implications extend to the breach of patient trust and the compromise of patient autonomy. The ethical medical practice relies heavily on the principles of informed consent and respect for patient’s autonomy—principles that are challenged by opaque AI systems that do not make their decision-making processes or inherent biases clear to patients or practitioners [77]. These potential legal and ethical failures underscore the necessity for rigorous oversight and transparent development processes for AI in healthcare, aiming to ensure that these technologies adhere to both existing legal frameworks and ethical standards of practice.
@@
Resource Misallocation: The deployment of biased AI in healthcare settings can lead to resource misallocation, a critical issue that impacts both the efficiency and fairness of medical services. Biased algorithms may misdirect resources by prioritizing certain groups over others based on flawed data inputs or biased training procedures. For example, a study by Obermeyer demonstrated how an algorithm used for managing healthcare resources inadvertently favored healthier white patients over sicker black patients due to biased data inputs that did not accurately reflect patient needs [12]. This misallocation can exacerbate existing healthcare disparities by diverting necessary medical attention and resources away from those who are most in need. As Chen pointed out, such disparities are not just a matter of clinical outcomes but are deeply tied to social and economic inequalities that AI tools can inadvertently perpetuate [13]. Furthermore, biased AI can influence the allocation of resources within healthcare facilities, potentially resulting in inefficiencies that strain healthcare systems. This includes misallocating medical staff, diagnostic tools, and hospital beds, which can degrade the quality of care delivered while increasing wait times and healthcare costs [4]. Resource misallocation also raises ethical questions about fairness and equity in healthcare provisioning. It challenges the ethical principle of justice, which demands that healthcare resources be distributed based on need rather than biased algorithms [81]. This ethical breach can lead to further mistrust and reluctance among underserved populations to engage with healthcare systems, perpetuating a cycle of disadvantage.
@@
Stifling Innovation: The presence of bias in AI systems used in healthcare not only affects the accuracy and fairness of medical services but can also stifle innovation. When AI models are made using biased data, they might not accurately reflect the different needs of the population. This could make these innovations less useful and applicable to a wider range of demographic groups. Liang talks about how biased training datasets in AI development make it harder for these systems to work with different types of data. This could stymie innovation by preventing it from solving larger or more complex health problems that affect many people. Moreover, reliance on biased AI can deter investment in developing new technologies that are inclusive and equitable. Potential investors may be cautious about funding projects that might not meet regulatory standards for fairness or could lead to public backlash — a concern highlighted by Corbett-Davies and Goel, who note the legal and social implications of deploying biased AI [83]. In addition, the perpetuation of bias in AI could solidify existing disparities in healthcare innovation. As Benjamin notes, if innovation is directed predominantly at solving problems specific to well-represented groups, less common but equally pressing issues in underrepresented groups are neglected, thereby limiting the scope and impact of technological advancements in healthcare [84]. The studies of Rajkomar [4], Hardt [85], and Howell [86] made things even more complicated by pointing out that biased algorithms can make it harder to find new treatments that work for everyone. This feeds a cycle of innovation that helps groups that are already overrepresented in data sets more than others. This restricted focus on innovation not only affects the equity of healthcare delivery but also limits the development of truly innovative, comprehensive healthcare solutions. These factors combined suggest that biased AI not only hinders the progression of medical technologies but also potentially locks the healthcare sector into a cycle of uneven innovation where only the needs of the majority are systematically addressed.
2.3 Fairness Metrics and Trade-offs
Biases in AI healthcare systems, as discussed in the previous sections, can significantly impact both the quality and equity of healthcare delivery. To address these challenges, fairness metrics provide a structured framework for evaluating and mitigating disparities in AI model predictions. Metrics such as Statistical Parity [87], Equal Opportunity [88], and Equalized Odds [89] serve as foundational tools for detecting bias. However, these metrics often reflect distinct fairness priorities, leading to inherent trade-offs when applied simultaneously [90]. For example, optimizing Statistical Parity to ensure equal outcomes across groups may conflict with Equal Opportunity, which focuses on ensuring equal access to positive outcomes for individuals who qualify. These conflicts underscore the complexity of achieving fairness in healthcare AI.
Furthermore, tensions arise not only between fairness metrics but also between different dimensions of fairness, such as group fairness and individual fairness. While group fairness seeks equitable treatment across demographic groups, individual fairness emphasizes similar outcomes for individuals with comparable features. These trade-offs become even more challenging in healthcare scenarios where fairness must often be balanced with performance, such as ensuring accurate predictions while maintaining equity. By navigating these conflicts, we aim to design AI systems that uphold fairness as a core principle in healthcare. To provide a clear understanding, we introduce these three key fairness metrics along with their mathematical foundations.
Statistical Parity (SP): Statistical Parity [87] ensures that the probability of a positive predicted outcome is equal across different groups defined by a sensitive attribute. Mathematically, it is defined as:
$SP=P(\hat{Y}=1\mid S=s_{1})-P(\hat{Y}=1\mid S=s_{2})$ (1)
where $\hat{Y}$ is the predicted outcome and $S$ is the sensitive attribute. Statistical Parity is satisfied when SP= 0, indicating no disparity between the groups.
Equal Opportunity (EO): Equal Opportunity [88] ensures that individuals who are likely to achieve the advantaged outcome (e.g., repaying a loan, being admitted to college) have an equal chance of receiving the positive prediction, regardless of their sensitive attribute (e.g., race, gender). Mathematically, it is defined as:
$EO=P(\hat{Y}=1\mid S=s_{1},Y=1)-P(\hat{Y}=1\mid S=s_{2},Y=1)$ (2)
where $\hat{Y}$ is the binary prediction, $S$ is the sensitive attribute, and $Y=1$ represents the advantaged outcome. Equal Opportunity is satisfied when $EO=0$ indicates no disparity in predictions for the advantaged outcome across sensitive groups.
Equalized Odds (Eq.Odds): Equalized Odds [89] ensures that for each ground truth label (Y = y), the predicted outcome $\hat{Y}$ is equally likely across groups defined by the sensitive attribute (S). Mathematically:
[…]
where $\hat{Y}$ is the predicted outcome, $S$ is the sensitive attribute, and $Y$ is the ground truth label. Equalized Odds is satisfied when $Eq.Odds=0$, indicating no disparity across groups for all ground truth labels.
These group fairness metrics Statistical Parity, Equal Opportunity, and Equalized Odds reflect fundamentally different fairness priorities, leading to inherent trade-offs when attempting to satisfy them simultaneously. Statistical Parity requires predictions to be independent of sensitive attributes, ensuring equal positive outcome rates across groups, regardless of differences in their underlying characteristics, which can sometimes lead to disparities in how individuals within a group are treated. Consider a healthcare scenario where males have a higher prevalence of heart disease compared to females, enforcing Statistical Parity would mean predicting the same proportion of males and females as being at risk. However, this might conflict with Equal Opportunity, which requires equal true positive rates across groups. Achieving Statistical Parity in this scenario may lower the true positive rate for females, as increasing their positive predictions could include many individuals without heart disease. This highlights how prioritizing Statistical Parity can undermine Equal Opportunity by creating disparities in prediction accuracy within groups.
To achieve Equalized Odds, a model must ensure that true positive and false positive rates are the same across groups, which often requires adjusting prediction thresholds differently for each group. For example, if females are less likely to have heart disease compared to males, the model might lower the prediction threshold for females to classify more of them as “at risk” increasing the true positive rate for females. At the same time, it might raise the threshold for males to reduce false positives. While these adjustments equalize predictive performance across groups, they disrupt Statistical Parity by creating unequal proportions of positive predictions for males and females, reflecting the differences in the underlying likelihood of heart disease between the groups. This trade-off shows how achieving Equalized Odds can violate Statistical Parity by altering the balance of positive predictions across groups.
The trade-offs between these fairness metrics are formalized in the Impossibility Theorem  [90], which demonstrates that it is often impossible to satisfy all fairness criteria simultaneously. For instance, achieving Statistical Parity might require selecting more from a group with a lower base rate, sacrificing Equalized Odds or Equal Opportunity, while prioritizing Equal Opportunity or Equalized Odds might allow selection rates to differ across groups, violating Statistical Parity. These inherent conflicts require careful consideration when deciding which fairness metric to apply. Ultimately, the choice of fairness metric should depend on the specific context and the ethical or practical priorities of the application. For instance, in critical healthcare scenarios like diagnosing heart disease, Equal Opportunity or Equalized Odds might be prioritized to ensure equal access to accurate treatment and minimize disparities in predictive performance. On the other hand, Statistical Parity could be more relevant in preventive care programs, where equitable outreach and representation are paramount goals [91].
The conflict between fairness metrics can also be seen in the difference between individual and group fairness, which focuses on fairness at different levels. Group fairness [92] aims to ensure equitable treatment across different subgroups defined by sensitive attributes, i.e., group fairness aims to ensure that the outputs of ML algorithms are independent of the sensitive attribute. In contrast, individual fairness [163] enforces fairness at a finer granularity by ensuring that similar individuals receive similar predictions based on their input features. Unfortunately, the simultaneous achievement of both maxima is practically impossible due to the trade-offs involved. Fig 1 illustrates this trade-off in the context of organ allocation, where gender is the sensitive attribute, and medical urgency determines priority. In the performance-driven result (left side of the Fig 1), organs are allocated solely based on urgency and likelihood of survival, prioritizing patients $P_{1}$, $P_{2}$, $P_{3}$, and $P_{4}$ due to their higher need and better survival prospects, while rejecting $P_{5}$, $P_{6}$, $P_{7}$, and $P_{8}$ because of their lower urgency or expected benefit. This approach maximizes utility by allocating resources to those expected to achieve the greatest overall benefit, such as improved survival rates and long-term health outcomes. However, it may inadvertently perpetuate systemic disparities between subgroups. The group fairness-driven result (right side of the Fig 1) adjusts the decision boundary to ensure equal approval rates for males and females, resulting in females $P_{7}$ and $P_{8}$ being approved despite their lower urgency, while males $P_{3}$ and $P_{4}$, with higher urgency, are denied. While this adjustment addresses equity at the group level, it compromises individual fairness, as highlighted by the purple rectangle, where individuals near the decision boundary disproportionately bear the group fair loss of the entire subgroup.
Building on the discussion of trade-offs between individual and group fairness, the tension between fairness and performance further illustrates these challenges. In contexts such as organ allocation, performance-driven approaches prioritize outcomes for individuals with the highest urgency or expected benefit. However, enforcing group fairness constraints such as ensuring equal approval rates across subgroups defined by sensitive attributes like gender can lead to a reduction in overall performance. For example, in the organ allocation scenario, ensuring that approval rates are equal for males and females might result in allocating organs to individuals like $P_{7}$ and $P_{8}$ (females, lower urgency) while denying organs to $P_{3}$ and $P_{4}$ (males, higher urgency). This adjustment ensures equity at the group level but comes at the cost of reduced overall utility, as fewer lives may be saved or long-term health outcomes may be compromised. This trade-off highlights the inherent difficulty of simultaneously achieving fairness and optimal performance, as prioritizing one objective often necessitates sacrifices in the other.
Fig 1: Illustration of the trade-off between performance-driven result and group fairness-driven result in the context of organ allocation, where gender is the sensitive attribute.